OpenStudy (anonymous):

-3,9,-27,81,... what is the pattern and the next 2 numbers??

OpenStudy (anonymous):

If you break them up into their prime numbers the pattern is easy to see.

OpenStudy (anonymous):

-3 = -1 * 3 9 = -1 * -1 * 3 * 3 -27 = -1 * -1 * -1 * 3 * 3 * 3 81 = -1 * -1 * -1 * -1 * 3 * 3 * 3 * 3